OpenStudy (anonymous):

Is the answer $85? Lawson will pay $612 to three different credit cards: Expert Card, Passport, and US Speedy. He will pay five times as much to Expert Card as US Speedy since it has a higher balance. He also expects the Passport payment to be only 20% of the total of the other two payments. How much will Lawson pay towards the Passport credit card balance?

OpenStudy (jdoe0001):

how did you get 85 anyway?

OpenStudy (anonymous):

No The Answer is $102:let the lowest price equal X and all the others be relative to X US is the lowest price so US=X Expert card is 5 times greater than US so Expert card is equal to 5X Passport is one fifth of 6X which is 1.2X You are solving for Passport so $612/7.2X(total of X's) = $85 $85x1.2=$102

OpenStudy (jdoe0001):

$$ \bf \text{a = us speedy amount}\\ 612 = a + 5a + 0.2(a+5a)\\ 612 = 6a + 0.2a + a \implies 612 = 7.2a\\ a = 85\\ 0.2(a+5a) \implies 0.2(85+5\times 85) \implies \large 102 $$

OpenStudy (jdoe0001):

:)